OpenStudy (anonymous):

A sequence... 5+3+2=151022 9+2+4=183652 8+6+3=482466 5+4+5=202541 then 9+5+5=??

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

I'd love to help, but don't see the pattern we need in order to evaluate 9+5+5. Are you positive that you've copied all of the info that came with this problem? Have you an example from class, notes, textbook or online to help you get started? I have very limited time this morning, unfortunately, since i have visitors coming.

OpenStudy (anonymous):

thats all the information i got... and maybe you got my message about sequence of first four digits so there should be some trick in last two...

OpenStudy (shrutipande9):

i found d trick of the 1st four nos..last 2 not yet..:/

OpenStudy (anonymous):

thats where i am stuck..

OpenStudy (mathmale):

Happy for this discussion, and wish I could be of help, but I really must get ready now for my visitors. Best wishes to you.

OpenStudy (anonymous):

thanks for your wishes... :) @mathmale

OpenStudy (anonymous):

143547 :)

OpenStudy (anonymous):

explain!!

OpenStudy (anonymous):

Answer : 143547 How? : 5+3+2 = 151022, means.. 5 x 3 = 15 5 x 2 = 10 3 + 2 = 5, 5 x 5 = 25, 25 - 3 = 22 and so..151022 7+2+5 =? 7 x 2 = 14 7 x 5 = 35 2 + 5 = 7, 7 x 7 = 49, 49 - 2 = 47 Answer : 143547 :)

OpenStudy (anonymous):

thanks...!!

OpenStudy (anonymous):

np :)

OpenStudy (raden):

5+3+2=(5x3)(5x2)(5x3+5x2 - 3)=151022 9+2+4=(9x2)(9x4)(9x2+9x4 - 2)=183652 8+6+3=(8x6)(8x3)(8x6+8x3 - 6)=482466 5+4+5=(5x4)(5x5)(5x4+5x5 - 4)=202541 then 9+5+5=?? 9+5+5=(9x5)(9x5)(9x5+9x5 - 5)=454585